OpenStudy (anonymous):

What is the solution to the equation? 5(8 + 7) = 5 • x + 5 • 7 A. x = 1 B. x = 5 C. x = 7 D. x = 8

OpenStudy (hysusonic):

what do you think we do first?

OpenStudy (anonymous):

Add 8 and 7 together then multiply the answer by 5 first.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so i got 20

OpenStudy (anonymous):

so what do i do next

OpenStudy (anonymous):

No that's not correct. add 8 and 7 together then multiply the answer by 5

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

75

OpenStudy (anonymous):

ok i got it it is a x=1

OpenStudy (anonymous):

am i right

OpenStudy (ghostgate):

Hold on, let me check.

OpenStudy (ghostgate):

It would be x = 8. Here's why: 5(8 + 7) = 5 * x + 5 x 7 5(15) = 5x + 35 75 = 5x + 35 40 = 5x 8 = x or x = 8

OpenStudy (ghostgate):

Because 5 x 8 = 40, plus 7 x 8 which is 35 = 75. Which 5(8 + 7) = 75. So 75 = 75.

OpenStudy (anonymous):

oh ok thank you very much

OpenStudy (ghostgate):

You're welcome! ;)

OpenStudy (anonymous):

i am good at math but this one was hard i have one more problem if you can help

OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 6 94 = 6 + 6 4 =

OpenStudy (ghostgate):

Is it 6 x 94 or what?

OpenStudy (ghostgate):

@taylor_152d can you please explain what it is? The spaces don't help identify the equation.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it says 6x94=6__?_+6x4=---?-

OpenStudy (ghostgate):

Well, 6 x 94 = 564. Hmmm, as for 6 x 4 = 24 + _ = ___. Hmmm.

OpenStudy (ghostgate):

Does 24 + _ = 564? If so, that can help narrow the problem.